OpenStudy (anonymous):

How much is 19.5% of 600?

OpenStudy (anonymous):

If you take 12% of 750, what do you get?

OpenStudy (dragons1415):

117

OpenStudy (nirmalnema):

x/600 * 100 = 19.5 x/6=19.5 x = 19.5*6 x=117

OpenStudy (dragons1415):

for the first question^^

OpenStudy (anonymous):

kk

OpenStudy (dragons1415):

second answer is 90 because 12% as a decimal is 0.12, multiply that by 750 and you get the answer of 90. Hope that helped :)

OpenStudy (anonymous):

thanks!

OpenStudy (dragons1415):

ur welcome :)

OpenStudy (anonymous):

if I need more help, may you help me?

OpenStudy (anonymous):

How much is 8% of 1,249?

OpenStudy (dragons1415):

yeah i'll help

OpenStudy (dragons1415):

the answer is 99.92 :)

OpenStudy (dragons1415):

any others? :P

OpenStudy (anonymous):

Find the rate in the following problem. (Round to three decimal places before changing to percent.) What percent of 142 is 317?